Sarah had been contributing $300 pre-tax per month to a retirement account that pays 2.16% interest compounded monthly. After 10 years, she needs to withdraw 25% of the money from her account. If the early withdrawal penalty is 10% of the amount withdrawn, how much will she have to pay? A. $1003.59 B. $40,143.57 C. 4,014.36 D. 10,035.89
1
Expert's answer
2022-01-06T06:46:01-0500

Solution:

Total payments="\$300\times12\times10=\$36,000"


Future value, F"= A\times\frac{(1+rate)^{nper }- 1}{rate }"


"rate = (1+\frac{r}{n})^{\frac{n}{p}}-1"


"n\ per = p \times t"


Where;

  • r = nominal annual interest rate (decimal)
  • n = number of compounding periods per year
  • p = number of payment periods per year
  • rate = rate per payment period
  • n per = total number of payment periods
  • A = an amount added to the principal at the end of each payment period


Future value, F"= 300\times\frac{(1.0018)^{120 }- 1}{0.0018}=40143.57"


Withdrawing "25\%" and incurring a "10\%" withdrawal fee, she will pay:


"0.25\times0.1\times40143.5=\$1003.59"

So, option A. is correct.